OpenStudy (anonymous):

Calculus Help! Use L'Hospital's Rule to show that "Limit as X approaches infinity that (1+1/x)*x=e.

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}\left( 1+1/x \right)^x=e\] That is my equation but I don't know how I use L"hospital's Rule here...